Justin is interested in buying a digital phone. He visited 12 stores at random and recorded the price of the particular phone he wants. The sample of prices had a mean of 342.25 and a standard deviation of 9.67.

(a) What t-score should be used for a 95% confidence interval for the mean, ?, of the distribution?
t* =

(b) Calculate a 95% confidence interval for the mean price of this model of digital phone:
(Enter the smaller value in the left answer box.)
to
Answer #1

Solution :

Given that,

\bar x = 342.25

s =9.67

n =12

Degrees of freedom = df = n - 1 =12 - 1 = 11

At 95% confidence level the t is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

  \alpha/ 2= 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,11 = 2.201 ( using student t table)

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.201 * (9.67 / \sqrt 12)

= 6.1441

The 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

342.25 - 6.1441 < \mu < 342.25+ 6.1441

336.1059 < \mu < 348.3941

(336.1059 to  348.3941 )